Beth is 4 years older than 3 times her cousin's age. If her couisn is 6, how old is Beth?

Found 2 solutions by JulietG, addingup:
Answer by JulietG(1812)

B = 3C + 4
If C = 6, then B = (3*6) + 4
If the cousin is 6, then Beth is 18+4, or 22

Answer by addingup(3677)

B = 3C+4 and C= 6
B = 3(6)+4 multiply 3*6 and add 4. Use your calculator. Or do it in your head, it's 22